# Optimal Computation in Anonymous Dynamic Networks

### v6

@@ Abstract
We give a simple characterization of which functions can be computed deterministically by anonymous processes in dynamic networks, depending on the number of leaders in the network. In addition, we provide efficient distributed algorithms for computing all such functions assuming minimal or no knowledge about the network. Each of our algorithms comes in two versions: one that terminates with the correct output and a faster one that stabilizes on the correct output without explicit termination. Notably, these are the first deterministic algorithms whose running times scale linearly with both the number of processes and a parameter of the network which we call dynamic disconnectivity (meaning that our dynamic networks do not necessarily have to be connected at all times). We also provide matching lower bounds, showing that all our algorithms are asymptotically optimal for any fixed number of leaders.
@@ 1.2 State of the Art
In convergent algorithms, the consensus is not reached in finite time, but each process’s local value asymptotically converges to the average. A state-of-the-art $\epsilon$-convergent algorithm based on Metropolis rules with a running time of $O\left(\tau n^{3}\log(1/{\epsilon})\right)$ communication rounds is given in [44]. However, this algorithm rests on the assumption that the degree of each process in the network has a known upper bound. [^4] More generally, averaging algorithms based on Metropolis rules cannot be applied to our network model, because they require all processes to know their outdegree at every round prior to sending their messages. The running time can be improved to $O\left(n^{2}\log(1/\epsilon)\right)$ communication rounds if the network is always connected (i.e., $\tau=1$) and may only change every three rounds (as opposed to every round).
@@ 1.3 Our Contributions
Problem class Leaders Terminating Assumptions Running time Lower bound
Frequency-based (e.g., Average Consensus) $\ell=0$ $\tau(2n-2)+1$ $\tau(2n-6)$
$\tau$ and $N\geq n$ known $\tau(n+N-2)+1$ $\tau(2n-4)$
$d$ known $\tau(n-1)+d+1\leq dn+1$
[…]
$\ell$ and $\tau$ known $\tau((\ell^{2}+\ell+1)(n-1)+1)$ $\tau(2n-\ell-3)$
@@
For each of the aforementioned functions, we give two algorithms: a terminating version, where all processes are required to commit on their output and never change it, and a slightly more efficient stabilizing version, where processes are allowed to modify their outputs, provided that they eventually stabilize on the correct output. The running times of our algorithms are summarized in Table 1.
@@
  • Assumptions on the network. Unlike most previous solutions, our algorithms work in any dynamic network that has a finite dynamic disconnectivity.
@@
Previous versions. This work is an extension of two preliminary conference papers that appeared at FOCS 2022 [23] and DISC 2023 [25], respectively. In preparing the current version, we have reworked several sections, included all missing proofs, and provided a more accurate analysis of the performance of our algorithms. In addition, the upper bound in terms of the dynamic diameter $d$ for leaderless networks is a new contribution (see Table 1). We also added a comparisons between history trees and previous structures and concepts related to anonymous and dynamic networks (see Section 3.3).
@@ 4.2 Stabilizing Algorithm for Leaderless Networks
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and stabilizes in at most $\tau(2n-2)+1$ rounds, assuming no knowledge of $\tau$ or $n$.
@@
Our local algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, output “Unknown”. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, output
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, at any round $t^{\prime}\geq\tau(2n-2)+1$, Condition 2 of Lemma 4.2 is met (since $c\geq 1$), and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $2\tau n$ rounds. ∎
@@ 4.4 Stabilizing Algorithm for Networks with Leaders
There is an algorithm that computes $F_{IM}$ in all $\tau$-union-connected anonymous networks with $\ell\geq 1$ leaders and stabilizes in at most $\tau(2n-3)+1$ rounds, assuming that $\ell$ is known to all processes, but assuming no knowledge of $\tau$ or $n$.
@@
The algorithm proceeds as in Theorem 4.4, with two differences. First, instead of returning “Unknown”, the new algorithm returns $\ell$. Second, when the fractions $\beta_{1}$, $\beta_{2}$, …, $\beta_{k^{\prime}}$ have been computed, as well as their sum $\beta$, we perform the following additional steps. Let $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$, and let $\{v_{j_{1}},v_{j_{2}},\dots,v_{j_{l}}\}\subseteq L_{0}$ be the set of nodes in $L_{0}$ representing leader processes, [^22] In general we have $l\leq\ell$, because some nodes of $L_{0}$ may represent more than one leader. i.e., such that $\text{label}(v_{j_{i}})$ has the leader flag set for all $1\leq i\leq l$. Compute $\beta^{\prime}=\sum_{i=1}^{l}\beta_{j_{i}}$ and $\gamma_{i}=\ell\beta_{i}/\beta^{\prime}$ for all $1\leq i\leq k^{\prime}$, and output
@@
If there are both leader and non-leader processes in the system, there are $c\geq 2$ distinct inputs, and therefore the stabilization time is at most $\tau(2n-3)+1$ rounds, as desired. In the special case where all processes are leaders, we have $\ell=n$, which is the algorithm’s output at every round since round 00 (recall that, by design, the algorithm returns $\ell$ instead of “Unknown”). Thus, the stabilization time is $\tau(2n-3)+1$ rounds in any case. ∎
@@ 5.2 Main Algorithm
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let ϑsuperscriptitalic-$\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@ 8   For $x:=\ell$ downto $1$
9      Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$      # see Listing 3 in Section 5.3
@@ 32   Else return $(n^{\prime},t)$
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let ϑsuperscriptitalic-$\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:

### v7

@@ Abstract
We give a simple characterization of the functions that can be computed deterministically by anonymous processes in dynamic networks, depending on the number of leaders in the network. In addition, we provide efficient distributed algorithms for computing all such functions assuming minimal or no knowledge about the network. Each of our algorithms comes in two versions: one that terminates with the correct output and a faster one that stabilizes on the correct output without explicit termination. Notably, these are the first deterministic algorithms whose running times scale linearly with both the number of processes and a parameter of the network which we call dynamic disconnectivity (meaning that our dynamic networks do not necessarily have to be connected at all times). We also provide matching lower bounds, showing that all our algorithms are asymptotically optimal for any fixed number of leaders.
@@ 1.2 State of the Art
In convergent algorithms, the consensus is not reached in finite time, but each process’ local value asymptotically converges to the average. A state-of-the-art $\epsilon$-convergent algorithm based on Metropolis rules with a running time of $O\left(\tau n^{3}\log(1/{\epsilon})\right)$ communication rounds is given in [44]. However, this algorithm rests on the assumption that the degree of each process in the network has a known upper bound. [^4] More generally, averaging algorithms based on Metropolis rules cannot be applied to our network model, because they require all processes to know their outdegree at every round prior to sending their messages. The running time can be improved to $O\left(n^{2}\log(1/\epsilon)\right)$ communication rounds if the network is always connected (i.e., $\tau=1$) and may only change every three rounds (as opposed to every round).
@@ 1.3 Our Contributions
For each of the aforementioned functions, we give two algorithms: a terminating version, where all processes are required to commit to their output and never change it, and a slightly more efficient stabilizing version, where processes are allowed to modify their outputs, provided that they eventually stabilize on the correct output. The running times of our algorithms are summarized in Table 1.
@@
  • Assumptions on the network. Unlike most previous solutions, our algorithms work in any dynamic network that has a finite dynamic disconnectivity (i.e., a finite dynamic diameter).
@@
Previous versions. This work is an extension of two preliminary conference papers that appeared at FOCS 2022 [23] and DISC 2023 [25], respectively. In preparing the current version, we have reworked several sections, included all missing proofs, and provided a more accurate analysis of the performance of our algorithms. In addition, the upper bound in terms of the dynamic diameter $d$ for leaderless networks is a new contribution (see Table 1). We also added comparisons between history trees and previous structures and concepts related to anonymous and dynamic networks (see Section 3.3).
@@ 4.2 Stabilizing Algorithm for Leaderless Networks
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and stabilizes in at most $\tau(2n-2)$ rounds, assuming no knowledge of $\tau$ or $n$.
@@
Our local algorithm is as follows. Each process $p$ runs Listing 1 on its own view $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, then $p$ outputs $\{(\lambda(p),1)\}$, where $\lambda(p)$ is its own input. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, $p$ outputs
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, if not all processes have the same input (hence $n\geq c\geq 2$), then at any round $t^{\prime}\geq\tau(2n-2)\geq\tau(2n-3)+1\geq\tau(2n-c-1)+1$, Condition 2 of Lemma 4.2 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $\tau(2n-2)$ rounds.
In the special case where all processes have the same input, then the default output $\{(\lambda(p),1)\}$ is correct since round 00. Also, whenever $S$ is a homogeneous system of $k-1$ independent linear equations, the algorithm still returns the correct output $\{(\lambda(p),1)\}$, because all nodes in level $L_{0}$ have the same label $\lambda(p)$. Hence, in this particular case the output is always correct. ∎
@@ 4.4 Stabilizing Algorithm for Networks with Leaders
There is an algorithm that computes $F_{IM}$ in all $\tau$-union-connected anonymous networks with $\ell\geq 1$ leaders and stabilizes in at most $\tau(2n-2)$ rounds, assuming that $\ell$ is known to all processes, but assuming no knowledge of $\tau$ or $n$.
@@
The algorithm proceeds as in Theorem 4.4, with two differences. First, the default output of a process $p$, instead of being $\{(\lambda(p),1)\}$, is now $\{(\lambda(p),\ell)\}$ (recall that the function $F_{IM}$ does not return frequencies, but numbers of processes). Second, when the fractions $\beta_{1}$, $\beta_{2}$, …, $\beta_{k^{\prime}}$ have been computed, as well as their sum $\beta$, the following additional steps are performed. Let $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$, and let $\{v_{j_{1}},v_{j_{2}},\dots,v_{j_{l}}\}\subseteq L_{0}$ be the set of nodes in $L_{0}$ representing leader processes, [^22] In general we have $l\leq\ell$, because some nodes of $L_{0}$ may represent more than one leader. i.e., such that $\text{label}(v_{j_{i}})$ has the leader flag set for all $1\leq i\leq l$. Compute $\beta^{\prime}=\sum_{i=1}^{l}\beta_{j_{i}}$ and $\gamma_{i}=\ell\beta_{i}/\beta^{\prime}$ for all $1\leq i\leq k^{\prime}$, and output
@@
If there are both leader and non-leader processes in the system, there are $c\geq 2$ distinct inputs, and therefore the stabilization time is at most $\tau(2n-2)\geq\tau(2n-c-1)+1$ rounds, as desired. In the special case where all processes are leaders, we have $\ell=n$, and the algorithm returns the correct output at every round since round 00. Thus, the stabilization time is $\tau(2n-2)$ rounds in any case. ∎
@@ 5.2 Main Algorithm
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let $\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@ 8 For $x:=\ell$ downto $1$
9 Assign $(n^{\prime},t):=\text{ApproxCount}(\mathcal{V},s,x,\ell)$ see Listing 3 in Section 5.3
@@ 32 Else return $(n^{\prime},t)$
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let $\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then: